OpenStudy (anonymous):

These are all T/F Questions. 1. If the rnage of a function consists of just one number, then its domain also consists of one number(I have false for this) 2. If the graph of y = f(x) has an x-intercept at x=a, then the gaph of y=f(x+h) has an x intercept at x = a-h(I have false) 3.The natural domain of the tangent function is the set of all real numbers(I have false)

OpenStudy (zarkon):

if \(y=f(x)\) has an \(x\)-intercept at \(a\) then \(f(a)=0\) then evaluating \(f(x+h)\) at \(a-h\) we get \(f(a-h+h)=f(a+0)=f(a)=0\) looks like it is an \(x\)-intercept